OpenStudy (anonymous):

A single die is rolled. Find the probability of rolling an odd number or a number less than 4. The prob. of rolling an odd number is 3/6. The prob. of rolling a number less than four is 3/6. but what do i do after that? Add them?

OpenStudy (anonymous):

An odd # or a number less than 4 is satisfied with the following rolls: 1, 2, 3, or 5 There are 6 possible rolls. Your P is : (# of successful rolls)/(total # of rolls) = 4/6 = 2/3

OpenStudy (anonymous):

All good now, @ello_bello ?

OpenStudy (anonymous):

yeah that is very helpful.

OpenStudy (anonymous):

uw! Good luck to you in all of your studies and thx for the recognition! @ello_bello